# Annealing Approximation in Master-Node Network Model

### v1

@@ 6 System with absorbing state $m^{\ast}=0$ and $m^{\ast}=1$ – Case III
The stability analysis is summarized in figure 3, while the technical details are found in Appendix D.
@@
$\left\{\begin{array}[]{ll}\Omega_{0}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L\geq 0\text{ and }R\leq 0\right\}\setminus\{(0,0)\}\\ \Omega_{1}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L\leq 0\text{ and }R\geq 0\right\}\setminus\{(0,0)\}\\ \Omega_{\tilde{m}}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L<0\text{ and }R<0\right\}\\ \Omega_{01}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L>0\text{ and }R>0\right\}\end{array}\right.,$ (45)
with $\Omega_{0}^{(01)}\cup\Omega_{1}^{(01)}\cup\Omega_{\tilde{m}}^{(01)}\cup\Omega_{01}^{(01)}=\mathbb{R}^{2}\setminus\{(0,0)\}$. As shown in figure 3, the regions $\Omega_{0}^{(01)}$, $\Omega_{1}^{(01)}$ and $\Omega_{\tilde{m}}^{(01)}$ contain a single stationary point, which are $m^{\ast}=0$, $m^{\ast}=1$ and $m^{\ast}=\tilde{m}^{(01)}$, respectively. On the other hand, the region $\Omega_{01}^{(01)}$ displays two stationary points, 00 and $1$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{01}^{(01)}$. In Appendix D, it is shown that $10<\tilde{m}^{(01)}<$ 1 in this region, and
@@
In other words, the point $m=\tilde{m}^{(01)}$, which is an unstable fixed point, is a separatrix between these two stationary points, 00 and $1$.
@@
The stationary behviour of the point $(0,0)$ should be examined separately, since the equation of motion is $\dot{m}=0$ in this point when $A^{(4)}=\overline{A^{(1)}}=0$. This means that no dynamics takes place at this particular point and the stationary value for $m(t)$ equals the initial value.
@@ 7 The role of the master node
Table 2: Transition rates $A^{(2)}$, $\overline{B^{1}}$, $\overline{B^{3}}$, $\overline{A^{3}}$, $B^{(2)}$ and $B^{(4)}$.
@@
Note that these four cases are not disjoint, but one can see the potential impact of the master node in the dynamics. To facilitate the analysis, let us consider two extremal cases. Firstly, if the master node is decoupled by letting $\Gamma_{i}=0$ for all sites in (3), it leads to
@@ 8 Discussion
Model Analysis and Key Findings
The master-node network model aims to provide a comprehensive framework for understanding opinion dynamics. The model can be used as a general framework in the spirit of other already proposed frameworks [17, 18, 19]. Here, we saw that mean field analysis show that a significant part of the parameter space can be understood through three distinct regimes of absorbing states. Indeed, by choosing the parameters properly, the model can represent several traditional effects from opinion mechanics, like contrarians or inflexible. Here, by employing annealing approximation and mean-field analysis, we have reduced the complex many-body problem to a tractable form while preserving essential features of the dynamics. The most significant theoretical advance lies in the complete characterization of phase diagrams for all three cases - systems converging to complete disagreement ($m^{*}=0$), complete consensus ($m^{*}=1$), or both states depending on initial conditions. The identification of tricritical points at $(L,R)=(0,0)$ reveals fundamental limitations in our ability to connect disordered and ordered phases smoothly.
Interpretation of R and L Parameters
The quantities $R$ and $L$ emerge as fundamental order parameters representing competing directional influences in the network. However, they do not correspond solely to movements towards one choice or another, as there are positive and negative terms in each of them. And yet, physically, a larger $R$ always corresponds to a right-oriented influence, combining the tendency to adopt right-neighbor opinions against resistance to such adoption. Conversely, $L$ quantifies left-oriented influence through similar mechanisms. Their ratio determines the system’s chiral character, with $R>L$ favoring right-biased consensus and vice versa. These parameters capture the inherent asymmetry in opinion propagation through the network. Interestingly, depending on those quantities, we observe a fixed point with coexistence of both opinions, like previous models with inflexibles [20, 21] or with asymmetric influences [galam2004dynamic].
Phase Transitions and Critical Behavior Let us analyze separately the scenarios with a single trivial stationary point (cases I and II) and the case where two trivial fixed points are displayed (case III). In the former, three regions, $\Omega_{trivial}$ (corresponding to $\Omega_{0}^{(0)}$ and $\Omega_{1}^{(1)}$ in cases I and II, respectively), $\Omega_{nontrivial}$ (domains $\Omega_{+}^{(0)}$ and $\Omega_{-}^{(1)}$ in cases I and II, in this order) and $\Omega_{2}$ (the regions $\Omega_{0+}^{(0)}$ and $\Omega_{1-}^{(1)}$ in cases I and II, respectively) can be identified. The transition from $\Omega_{trivial}$ to $\Omega_{nontrivial}$ is smooth in the sense of (22) and (35), while a discontinuity is observed in the frontier between $\Omega_{nontrivial}$ and $\Omega_{2}$. Therefore, the point $(0,0)$ can be regarded as a tricritical point. The region $\Omega_{2}$ has two fixed points, where the initial condition of the dynamics plays a major role in determining to which one the system converges; furthermore, the boundary that separates to which stationary point the systems moves toward to is given by an unstable stationary point (see (21) and (34)). Finally, the transition from $\Omega_{nontrivial}$ and $\Omega_{2}$ is smooth because it presents the behavior associated with the nontrivial stationary point, shared by these two regions. On the other hand, in case III, two trivial stationary points ($m^{\ast}=0$ and $m^{\ast}=1$) are present and generate the domains $\Omega_{0}^{(01)}$ and $\Omega_{1}^{(01)}$. This time, the boundaries of these two domains are closed (in the sense that the stationary value of $m(t)$ matches the value at its core), except for the origin $(0,0)$, where the stationary point equals the initial value. The region $\Omega_{01}^{(01)}$, which is the analogous phase of $\Omega_{2}$, displays a stationary point according to the initial condition.
Implications and Future Directions These results provide significant insights for modeling real-world social systems, particularly in understanding how centralized influences can alter collective behavior. The phase diagrams offer predictive power for consensus formation. Future work should investigate the model beyond mean-field approximation and explore connections to measurable social quantities in empirical networks. The framework also suggests principles for engineering consensus in organizational systems.
Conclusion
This work has established a comprehensive theoretical framework for analyzing absorbing-state transitions in opinion dynamics through a master-node network model. The annealing approximation approach has revealed several fundamental insights about collective behavior in social systems:
1) The model demonstrates how simple interaction rules can generate complex phase diagrams with both continuous and discontinuous transitions between disordered and ordered states. The characterization of three distinct regimes (single absorbing state at $m^{*}=0$, at $m^{*}=1$, and bistable systems) provides a complete taxonomy of possible consensus formation scenarios.
2) The chiral parameters $R$ and $L$ emerge as powerful descriptors of system behavior, quantifying the inherent asymmetry in opinion propagation through directional influences. Their ratio determines the ultimate consensus state and helps explain polarization phenomena.
3) The identification of tricritical points and the detailed analysis of transition lines offer practical tools for predicting when small changes in network parameters may lead to dramatic shifts in collective opinion.
@@ Conclusion
4) The master node’s role as a symmetry-breaking element suggests design principles for influencing real-world social systems, from organizational management to public opinion shaping.
@@
A Conclusion
This study presents a significant contribution to the understanding of opinion dynamics within complex systems through a master-node model under an annealing approximation framework. By introducing a chiral formulation of the master node, we were able to uncover a rich phase diagram that includes both continuous and discontinuous transitions between ordered and disordered states. These transitions are highly dependent on initial conditions, particularly in systems with two absorbing states: complete disagreement ($m^{*}=0$) and full consensus ($m^{*}=1$). The identification of tricritical points at $(L,R)=(0,0)$ further highlights the limitations in smoothly connecting disordered and ordered phases, reinforcing the idea that consensus formation is not always a gradual process and can exhibit abrupt shifts under certain conditions.
Ultimately, this line of inquiry contributes to a deeper understanding of how information, influence, and consensus evolve in complex adaptive systems — a question of growing importance in our increasingly interconnected world.
B Appendix – case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$
As mentioned in the main text, the stability analysis of the stationary points are based on the (one-dimensional) isocline method. This means that one investigates the sign of $F_{0}$ in the differential equation $\dot{m}=F_{0}(m)$, which means that the relative positions of the key points 00, $1$, $m_{-}^{(0)}$ and $m_{+}^{(0)}$ are helpful. Although $m=1$ is not a stationary point, it is included in the analysis since the meaningful interval for $m$ is $[0,1]$. The analysis of $m_{-}^{(0)}$ and $m_{+}^{(0)}$ will be shown for reference even if they do not belong to this meaningful interval, though the physical outcome is restricted to the stability of these key points in $[0,1]$ only. The analysis is divided into some parts, as specified in table 3.

### v2

@@ 6 System with absorbing state $m^{\ast}=0$ and $m^{\ast}=1$: Case III
The stability analysis is summarized in figure 4, while the technical details are found in Appendix C.
@@
[…]
with $\Omega_{0}^{(01)}\cup\Omega_{1}^{(01)}\cup\Omega_{\tilde{m}}^{(01)}\cup\Omega_{01}^{(01)}=\mathbb{R}^{2}\setminus\{(0,0)\}$. As shown in figure 4, the regions $\Omega_{0}^{(01)}$, $\Omega_{1}^{(01)}$ and $\Omega_{\tilde{m}}^{(01)}$ contain a single stationary point, which are $m^{\ast}=0$, $m^{\ast}=1$ and $m^{\ast}=\tilde{m}^{(01)}$, respectively. On the other hand, the region $\Omega_{01}^{(01)}$ displays two stationary points, 0 and $1$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{01}^{(01)}$. In Appendix C, it is shown that $0<\tilde{m}^{(01)}<1$ in this region, and
@@
In other words, the point $m=\tilde{m}^{(01)}$, which is an unstable fixed point, is a separatrix between these two stationary points, 0 and $1$.
@@
The stationary behvior of the point $(0,0)$ should be examined separately, since the equation of motion is $\dot{m}=0$ in this point when $A^{(4)}=\overline{A^{(1)}}=0$. This means that no dynamics takes place at this particular point and the stationary value for $m(t)$ equals the initial value.
@@ 7 The role of the master node
Table 3: Transition rates $A^{(2)}$, $\overline{B^{(1)}}$, $\overline{B^{(3)}}$, $\overline{A^{(3)}}$, $B^{(2)}$ and $B^{(4)}$.
@@
Note that these four cases are not disjoint, but one can see the potential impact of the master node in the dynamics. To facilitate the analysis, let us consider two extremal cases. First, if the master node is decoupled by letting $\Gamma_{i}=0$ for all sites in (3), it leads to
@@ 8 Discussion
Model analysis and key findings. The master-node network model aims to provide a broad and unifying framework for the study of opinion dynamics, much like other general approaches in sociophysics [21, 22, 23]. The mean-field analysis presented here reveals that a large portion of the parameter space can be partitioned into three distinct absorbing-state regimes. By appropriately choosing the parameters, the model is flexible enough to reproduce classical mechanisms in opinion dynamics, such as the presence of contrarians or inflexibles. The adoption of the annealing approximation, together with mean-field techniques, transforms the otherwise complex many-body stochastic process into a tractable analytical framework while still retaining the essential ingredients of the dynamics. The main theoretical advance of this work lies in the complete characterization of phase diagrams for all three cases, covering convergence to complete disagreement ($m^{*}=0$), convergence to complete consensus ($m^{*}=1$), and bistability where the outcome depends on the initial condition of the system.
This study builds directly on the master-node network model previously introduced in [15, 16]. In those works, agents were arranged in a one-dimensional ring with nearest-neighbor interactions, and a master node exerted external influence over a quenched fraction $\rho$ of agents. The master node imposed its state with strength $r$, while $\Delta$ quantified the intrinsic bias of agents toward selfish ($s=0$) or cooperative ($s=1$) behavior. Across these formulations, the central object of interest was the absorbing-state phase transition, measured through the order parameter $I$ (or equivalently $m$), which describes the fraction of agents opposing the master node’s state.
Compared with [15], the present work provides a substantially more general formulation by introducing chirality via the parameters $L$ and $R$, which quantify left- and right-oriented asymmetries in influence. This extension greatly enriches the dynamical landscape, giving rise to both continuous and discontinuous transitions and unveiling the existence of tricritical organization around $(L,R)=(0,0)$. Another methodological advance is the adoption of the annealing approximation, in which the quenched disorder represented by $\rho$ is replaced by its mean value. This simplification makes the mean-field analysis analytically tractable and allows us to map out complete phase diagrams, classifying the system into three absorbing-state regimes. Consequently, the present model yields diagrams that are richer and more informative than those reported in [15], which primarily studied continuous transitions through Monte Carlo simulations and finite-size scaling techniques.
Relative to [16], which emphasized critical exponents obtained from extensive Monte Carlo simulations, the present study represents a complementary but equally important contribution. Although that earlier work demonstrated that the model does not belong to universality classes such as directed percolation or the voter model, here we adopt an analytical approach based on annealing approximation and mean-field theory. The inclusion of chirality through $L$ and $R$ allows for explicit equations of motion and the construction of complete phase diagrams. This broader formulation highlights how asymmetric interactions reshape consensus and polarization, generating dynamical regimes that could not be captured by the original three-parameter version $(r,\rho,\Delta)$.
Taken together, the three studies demonstrate continuity and progression. While all share a common structural foundation and focus on absorbing-state transitions in networks with a master node, the present contribution advances the field by generalizing the model to include chirality, developing an analytical framework that unifies continuous and discontinuous transitions, and uncovering tricritical organization. This places our results as a theoretical extension that consolidates and surpasses the findings of [15, 16].
Phase transitions and critical behavior. In order to clarify the critical behavior, let us examine separately the scenarios with a single trivial stationary point (Cases I and II) and the case with two trivial absorbing states (Case III). In the former, three dynamical regions can be identified: $\Omega_{\text{trivial}}$, where only the trivial fixed point exists ($m^{*}=0$ in Case I, $m^{*}=1$ in Case II); $\Omega_{\text{nontrivial}}$, where a new nontrivial stationary point emerges; and $\Omega_{2}$, where the system becomes bistable with coexistence of trivial and nontrivial solutions. The transition from $\Omega_{\text{trivial}}$ to $\Omega_{\text{nontrivial}}$ is continuous, in line with the behavior described in (22) and (35). By contrast, the transition from $\Omega_{\text{nontrivial}}$ to $\Omega_{2}$ is discontinuous and associated with a saddle-node bifurcation.
It is essential to emphasize that this saddle-node bifurcation take place along the boundary curves $R=\sqrt{4LA^{(4)}}$ (Case I) and $L=\sqrt{4R\overline{A^{(1)}}}$ (Case II). These boundaries separate the pink (monostable) and gray (bistable) regions in Fig. 2. At the boundary, a single nontrivial fixed point emerges and then splits into two distinct solutions inside the gray region: a stable node ($m_{+}$) and an unstable saddle ($m_{-}$). Crucially, the trivial solution ($m^{*}=0$ in Case I, or $m^{*}=1$ in Case II) remains present in both regions and is unaffected by the bifurcation.
The region $\Omega_{2}$ therefore contains two attractors, and the initial condition plays a decisive role in determining which stationary state the dynamics converge to. The unstable solution $m_{-}$ functions as a separatrix between the basins of attraction. The coexistence of two stable solutions ($m^{*}=0$ and $m_{+}$, or $m^{*}=1$ and $m_{+}$ depending on the case) reflects a first-order transition mediated by a saddle-node bifurcation.
In Case III, the situation differs in that both $m^{*}=0$ and $m^{*}=1$ are absorbing states. The corresponding domains, $\Omega_{0}^{(01)}$ and $\Omega_{1}^{(01)}$, are closed, in the sense that the stationary value of $m(t)$ matches the absorbing state. At the origin $(0,0)$ the dynamics are degenerate: The stationary point simply equals the initial condition. The mixed region $\Omega_{01}^{(01)}$ thus behaves analogously to $\Omega_{2}$ in Cases I and II, in that the system’s fate depends on initial conditions, but without the appearance of a discontinuous transition or a saddle-node bifurcation. Instead, $(0,0)$ plays a structural role by defining the polarization threshold $\tilde{m}^{(01)}=L/(R+L)$.
Altogether, these results provide a coherent picture: $(0,0)$ acts as a tricritical organizer, structuring the global phase diagram, while the actual saddle-node bifurcations that generate bistability occur along finite-LL/$R$ boundaries. This distinction clarifies the relationship between continuous and discontinuous transitions and shows how chirality governs the onset of bistability in the model. The parameters $L$ and $R$ thus determine not only the final state of consensus or polarization but also the very nature of the critical behavior—continuous, discontinuous, or tricritical.
By situating these results alongside previous work, we conclude that the inclusion of chirality and the use of annealing approximation not only generalize the master-node model but also reveal fundamentally new critical properties absent in the quenched-disorder formulations. This analytical framework deepens our understanding of how asymmetric influence can drive consensus, polarization, and bistability in social systems.
8.1 Interdisciplinary implications
Although our model is formulated in the language of mathematics and inspired by concepts from statistical mechanics, it aims to describe profoundly social phenomena: how asymmetric influences, network structures, and initial conditions shape collective opinion dynamics—ultimately leading to consensus, polarization, or abrupt collapse.
Abrupt transitions in opinion dynamics have sparked debate, as certain models yield qualitatively different predictions depending on how the thermodynamic limit is taken [24, 25, 26]. Our work resolves this ambiguity by providing a mathematically robust framework in which discontinuous transitions emerge naturally from the underlying dynamics of chiral asymmetry—rather than as artifacts of computational or limiting procedures. While grounded in the formalism of nonequilibrium statistical mechanics and phase transitions, the model’s central insights—particularly regarding asymmetric influence, tipping points, bistability, and consensus collapse—resonate strongly with long-standing questions in sociology, political science, and social psychology.
The notion of a master node exerting asymmetric, chiral influence aligns closely with sociological theories of opinion leadership and influencer dynamics in social networks [27, 28]. This conceptual parallel extends to Katz and Lazarsfeld’s seminal two-step flow of communication model [29], historically developed (with Edward Shils) to challenge the simplistic “hypodermic needle” theory of media effects [30]. Like our master node, opinion leaders in this framework act as pivotal intermediaries who channel, filter, and amplify external influence within social networks.
Directional bias—or chirality—in our model captures asymmetric influence, which arises naturally in contexts featuring clear hierarchies, such as those between opinion leaders and followers [27, 28], digital influencers [31], or when agents exhibit differential responsiveness to competing options [32]. This asymmetry is central to understanding how small imbalances can drive large-scale outcomes.
Our findings on discontinuous transitions and bistability echo Granovetter’s threshold models of collective behavior [33], which show that macro level shifts—such as protest cascades or market tipping points—depend critically on the distribution of individual thresholds, not merely their average. Similarly, in our framework, system-level outcomes hinge on the fine-grained structure of influence and initial conditions, rather than on aggregate statistical moments.
The identification of tricritical points and bistable regimes—where systems exhibit path dependence and resistance to perturbations—finds direct parallels in political science literature on policy feedback and institutional lock-in [34]. In both cases, once a system settles into an absorbing state, minor shocks are insufficient to reverse its trajectory; only structural interventions (in our model, changes in the parameters $L$ and $R$) can shift the equilibrium.
Moreover, the role of directional bias in driving polarization aligns with Sunstein’s account of echo chambers and group polarization [35]. In such settings, homophilic interactions amplify existing views not through manipulation, but via ordinary psychological mechanisms—social validation and reinforcement—mirroring the chiral feedback loops in our model.
Finally, the sensitivity of consensus outcomes to initial conditions—especially in bistable regimes—reflects well-established findings in social psychology on priming and framing effects [36]. Druckman demonstrates that framing influences preferences not automatically, but through mediating factors such as source credibility, individual predispositions, and deliberation. Our model incorporates this insight by treating the initial magnetization $m_{0}$ as a decisive variable that shapes the system’s long-term trajectory.
These mechanisms are especially relevant for modeling—and understanding—theoretical aspects of real-world social systems where a centralized source of influence (e.g., media outlets, political leaders, digital platforms, or parties) exerts asymmetric pressure on a distributed network of agents. In such contexts, collective outcomes often hinge critically on initial opinion configurations and the precise architecture of influence.
The framework also applies to scenarios where minor structural changes in influence networks trigger qualitative behavioral shifts, or where systems remain locked into specific opinion states even after the original source of influence is removed—phenomena described in the literature as path dependence in public policy [34] and persistent polarization in echo chambers [35].
Concrete examples include the following:
  • The spread of misinformation or political messaging on social media, where recommendation algorithms act as master nodes that asymmetrically promote content, potentially leading to abrupt consensus or deep polarization [35, 33];
  • Social mobilization and protest dynamics, where small groups of instigators succeed or fail based on the distribution of individual participation thresholds [33];
  • Institutional inertia in public policy, where systems remain entrenched in suboptimal equilibria not due to active opposition, but because of path-dependent lock-in [34];
  • Public health campaigns, whose effectiveness depends not only on message content but also on source credibility and how the population is initially framed [36, 37];
  • Market dynamics and technology adoption, which often exhibit abrupt transitions from niche to mainstream once a critical mass is reached [33, 38].
The interdisciplinary insights developed above—linking chiral asymmetry, initial-condition sensitivity, and abrupt transitions to sociological theories of influence, institutional lock-in, and echo chambers—find their most precise theoretical counterpart in the recent classification by Van Santen et al. [4], who established a foundational framework for opinion dynamics under centralized external influence. Their model distinguishes between symmetric and asymmetric interaction strategies, demonstrating how a master like agent can drive consensus either toward or against its own opinion, depending on network structure and clustering effects. This aligns directly with our core premise: that a single, asymmetrically connected node can fundamentally reshape collective outcomes by breaking the symmetry of peer interactions. In this broader context, our chiral master-node model both confirms and transcends Van Santen et al.’s paradigm [4]. While they identify macroscopic regimes—consensus, disorder, or anticonsensus — as continuous functions of an external field strength $E_{A}$, we reveal that under chiral asymmetry ($L\neq R$), the system undergoes qualitative phase transitions governed not by gradual drift, but by structural instabilities. The emergence of a saddle-node bifurcation along the boundary $L^{2}=4R\overline{A^{(1)}}$ (Case II) or $R^{2}=4LA^{(4)}$ (Case I) generates discontinuous jumps between stable states, a phenomenon absent in their continuous framework. Crucially, our work introduces two interrelated dimensions absent in Van Santen et al.’s classification [4]: (i) the existence of a tricritical point at $(L,R)=(0,0)$, which organizes the entire phase diagram by marking the termination of continuous transitions and the birth of discontinuous ones, and (ii) bistability arising from the geometric separation of stable fixed points ($m_{+}$ and $m_{-}$) along the boundary between monostable and multistable regions, as visualized in Fig. 2. These features are not mere quantitative extensions—they represent a qualitative shift in the nature of influence: from a tunable perturbation to a topological reorganization of the state space. Where Van Santen et al. [4] describe how social clustering amplifies or resists influence, we show how directional bias—encoded in the chiral parameters $L$ and RR—determines whether the system evolves smoothly toward consensus or fractures abruptly into polarization. In their model, the transition from disagreement to consensus is monotonic; in ours, it is geometrically constrained by the emergence of unstable separatrices, making outcomes exquisitely sensitive to initial conditions even in the absence of memory, inflexible agents, or heterogeneity. The point $(0,0)$ is not merely a parameter value—it is the apex of a bifurcation landscape where the very character of collective behavior changes. Thus, while Van Santen et al.[4] provide a powerful taxonomy of opinion regimes under centralized influence, our model unveils the underlying dynamical architecture: a phase space organized by chirality, punctuated by a tricritical singularity, and governed by irreversible, nonmonotonic transitions. We do not replace their framework—we embed it within a richer mathematical structure, revealing that the most profound societal shifts arise not from stronger signals, but from the geometry of asymmetry itself.
9 Conclusion
This study presents a significant contribution to the understanding of opinion dynamics within complex systems through a master node model under a mean-field annealing approximation framework. By introducing a chiral formulation of the master node, we were able to uncover a rich phase diagram that includes both continuous and discontinuous transitions between ordered (full consensus or complete disagreement) and disordered (coexistence of opinions) states, where the discontinuity is triggered by a saddle-node bifurcation. The chiral parameters $R$ and $L$ emerge as powerful descriptors of system behavior, quantifying the inherent asymmetry in opinion propagation through directional influences. They determine the ultimate consensus state and helps explain polarization phenomena. We also identified regions in the phase diagram where these transitions are highly dependent on initial conditions. The identification of tricritical points at $(L,R)=(0,0)$ further highlights the limitations in smoothly connecting disordered and ordered phases, reinforcing the idea that consensus formation is not always a gradual process and can exhibit abrupt shifts under certain conditions. The detailed analysis of transition lines offer practical tools for predicting when small changes in network parameters may lead to dramatic shifts in collective opinion.
Our starting point was a general model with several free parameters. Despite this complexity, its mean-field formulation, combined with an annealing approximation in the way the master node influences the spins, led to an unexpectedly simplified description governed by a specific combination of the original parameters, ultimately reducing them to four key quantities ($R$, $L$, $A^{(4)}$, and $\overline{A^{(1)}}$), which could be investigated analytically. While this constitutes our main contribution, it also represents a limitation, as the resulting framework deviates from the original model.
@@ 9 Conclusion
Ultimately, this line of inquiry contributes to a deeper understanding of how information, influence, and consensus evolve in complex adaptive systems — a question of growing importance in our increasingly interconnected world. The master node’s role as a symmetry-breaking element suggests design principles for influencing real-world social systems, from organizational management to public opinion shaping.
@@
10 Acknowledgements
FFF thanks CNPq for support, Grant No. 316664/2021-9.
A Appendix: case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$
As mentioned in the main text, the stability analysis of the stationary points are based on the (one-dimensional) isocline method. This means that one investigates the sign of $F_{0}$ in the differential equation $\dot{m}=F_{0}(m)$, which means that the relative positions of the key points 0, 11, $m_{-}^{(0)}$ and $m_{+}^{(0)}$ are helpful. Although $m=1$ is not a stationary point, it is included in the analysis since the meaningful interval for $m$ is $[0,1]$. The analysis of $m_{-}^{(0)}$ and $m_{+}^{(0)}$ will be shown for reference even if they do not belong to this meaningful interval, though the physical outcome is restricted to the stability of these key points in $[0,1]$ only. The analysis is divided into some parts, as specified in table 4.